# Benchmarking Gaussian and non‑Gaussian input states with a hybrid sampling platform

### v1

@@ Abstract
The original boson sampling paradigm—consisting of multiple single-photon input states, a large interferometer, and multi‑channel click detection—was originally proposed as a photonic route to quantum computational advantage. Its non‑Gaussian resources, essential for outperforming any classical system, are provided by single‑photon inputs and click detection. Yet the drive toward larger experiments has led to the replacement of experimentally demanding single‑photon sources with Gaussian states, thereby diminishing the available non‑Gaussianity—a critical quantum resource. As the community broadens its focus from the initial sampling task to possible real‑world applications, it becomes crucial to quantify the performance cost associated with reducing non‑Gaussian resources and to benchmark sampling platforms that employ different input states. To address this need, we introduce the Paderborn Quantum Sampler (PaQS), a hybrid platform capable of performing sampling experiments with eight Gaussian or non‑Gaussian input states in a 12‑mode interferometer within a single experimental run. This architecture enables direct, side‑by‑side benchmarking of distinct sampling regimes under otherwise identical conditions. By employing a semi-device-independent framework, offering certification that does not rely on prior knowledge of the interferometer or the input states, we verify that the observed data cannot be reproduced by any classical model—a prerequisite for demonstrating quantum advantage. Applying this framework, we observe clear performance gains arising from non‑Gaussian input states.
@@
The non‑Gaussian element in GBS systems arises from photon‑number‑resolved (PNR) detection (or in earlier implementations, click detection). Extending the computational capabilities of such systems is therefore most easily achieved by supplementing this resource, either through advanced operations within the interferometer [31, 34, 35, 36], or by driving the interferometer with more exotic non‑Gaussian input states, whose impact has been the subject of extensive investigation [37, 38, 39, 40, 41, 42, 43, 44]. Consequently, the next generation of sampling devices will require a conceptual redesign of both their system architecture and resource allocation.
@@
A corresponding rethink is necessary for benchmarking the performance of these newly conceived sampling systems. A commonly used verification technique is to quantify the closeness of generated data to that which is expected from theory, typically implemented using a metric known as the total variation distance [1]. However, this method does not serve as an ideal benchmarking tool for different sampling configurations, as it makes no direct statement about presence of quantumness in the system, particularly in the presence of system imperfections—a requirement to perform beyond classical limitations. Progressing the field therefore requires new certification tools capable of identifying genuinely non‑classical features—or quantumness—and experimental platforms that enable fair, side‑by‑side comparison of different architectures.
@@
In this work, we introduce the Paderborn Quantum Sampler (PaQS), an experimental platform designed to benchmark multiple sampling schemes within a single experimental run. To compare the performance associated with different input‑state resources—both Gaussian and non-Gaussian—as illustrated in Figure 1, we further develop a benchmarking framework based on normally ordered moments of the photon‑number operator [46, 47, 48, 49, 50, 51]. In contrast to previous approaches, this method certifies the presence of quantumness in the measured data— a necessary prerequisite for any claim of quantum computational advantage [52].
@@
Our results reveal stark differences between the generated GBS and SBS datasets. The quantumness of SBS data increases steadily with the mean photon number of the input states, whereas the GBS data displays strong non‑classical signatures at low mean photon numbers, but fails to maintain them as the brightness increases. This divergence underscores the fundamentally different behaviors arising from distinct input‑state resources and highlights the need for a conceptual shift when expanding sampling architectures to operate beyond their original task.
@@
Figure 1: Conceptual implementations for various sampling configurations. In all cases, the input fields enter from the bottom of the setup and are detected with PNR detectors upon exiting the interferometer at the top. The input states of the interferometer can be either eight SMSV states, heralded Fock states, or thermal states, corresponding to implementing GBS, SBS, and thermal boson sampling (TBS), respectively. To implement SBS, the second mode of each TMSV input state is detected with PNR detectors to herald the number of photons entering each interferometer port. To realize sampling with thermal states - thermal boson sampling (TBS) - the second mode of each TMSV state is discarded, i.e., traced out.
Experimental Setup
@@
Figure 2 presents a schematic of the PaQS system, divided into its constituent subsystems for clarity. The design of PaQS emphasizes integration: both the squeezed-light source (or parametric down-conversion source, PDC) and the programmable interferometer are realized on integrated photonic platforms. Notably, the integrated source, in conjunction with an electro-optic modulator and polarizing beamsplitter, enables dynamic switching between TMSV- and SMSV-state generation. This architecture allows multiple sampling configurations to be acquired during a single experimental run, thereby ensuring that results from the different sampling schemes are directly commensurable. Furthermore, the picosecond squeezed-light pulses generated by the waveguided source enables the implementation of intrinsic photon-number resolution (PNR) in the employed superconducting nanowire single-photon detectors (SNSPDs). The integrated interferometer offers convenient and complete programmability of the implemented unitary transformation, with an average insertion loss below 3 dB (see Supplementary Material). An overview of the key subsystems of PaQS is given in Figure 2, while detailed descriptions are provided in the Supplementary Material.
@@
A train of pulses is first selected from a recently developed Menlo laser system consisting of an ultra-low-noise stabilized frequency comb with a central wavelength of 1544 nm (Menlo Systems GmbH), a laser extension unit for providing control signals and laser amplification, and a frequency doubling stage to generate 772 nm light. This train of pulses is spectrally shaped to further optimize the pump for squeezed-light generation. The $\approx 2$ ps pulse duration squeezed-light pulses are produced in a single-pass, type-II periodically poled potassium titanyl phosphate (PPKTP) waveguide (AdvR Inc.), then spectrally filtered using a silicon window, to remove the pump filter, and a 2 nm bandpass filter, to remove the phase-matching side lobes. After filtering, the pulses are group-delay-compensated using a compensation crystal (CC) before passing through an electro-optic modulator (EOM2, QUBIG GmbH) and a polarizing beam splitter (PBS). By varying the voltage applied to EOM2, we can rapidly switch (rise/fall time of 4 ns) between generating two independent SMSV states and a TMSV state at the PBS output. This tunability enables the implementation of GBS [8], SBS [7], and TBS [53] configurations. One PBS output is directed into a multiplexed detection scheme comprising a splitting tree and eight SNSPDs (Single Quantum), while the other arm passes through an active time-to-space demultiplexer (QUBIG GmbH) and a path-length compensation system consisting of multiple path-length compensation units (PLCUs - Menlo Systems GmbH) that route the pulses into individual fibers and synchronize their arrival times.
The synchronized pulses subsequently interfere in a 12-mode, integrated, fully programmable interferometer (QuiX Quantum), of which eight channels are used. The photons present in each output channel are detected using eight SNSPDs (Single Quantum). Intrinsic photon-number detection up to three photons is realized on all detectors using a recently developed technique that determines photon number from the rise time of the SNSPD signal measured using a high resolution time tagger (Swabian Instruments GmbH) [54, 55]. Further experimental details are provided in the Supplementary Material.
System Characterization
@@
We summarize below the principal methods used to characterize and verify the operation of the PaQS subsystems. The corresponding results are shown in Figure 3, with further details provided in the Supplementary Material.
@@
Single spectro-temporal-mode operation of our type-II waveguide source is verified through second-order autocorrelation measurements [57]. The system operates in SBS mode while photon statistics are recorded for one output of the PBS as the mean photon number is varied by adjusting the pump power driving the nonlinear process. As shown in Figure 3 b, the second-order correlation function converges to $g^{(2)}=1.95\pm 0.03$ at higher power, corresponding to an expected Schmidt mode number $K=1.05\pm 0.03$ effective modes—confirming both spectro-temporal purity and suitability for sampling experiments. The increase in $g^{(2)}$ at low mean photon numbers originates from a residual SMSV contribution because of imperfect polarization separation at the PBS (see Supplementary Material).
Next, we assess the ability of PaQS to switch between SMSV and TMSV generation. Operating in the low‑gain regime, we record the coincidence counts $n_{\mathrm{C}}$ between the herald and signal detectors as the driving voltage of EOM2’s high-voltage amplifier is varied (Figure 3 c). The normalized coincidence signal varies nearly sinusoidally with a fitted visibility of $V=96.3\%\!\pm\!1.2\%$ (See Supplementary Material), demonstrating precise and continuous tunability between SBS (maximum coincidences) and GBS (minimum coincidences) configurations. The difference in number of normalized coincidence counts for the two displayed SBS configurations (at 0 V and $\approx$1.35 V) is due to the different spatial-mode profiles of the two polarization modes within the waveguide.
Finally, Figure 3 d illustrates the precision with which we can tune the optical path-length difference between input modes using the PLCUs in order to optimize the indistinguishability between different sources (further details in the Supplementary Material). Implementing the SBS configuration in the low‑gain regime, we set the interferometer to realize a 50:50 beam splitter between any pair of chosen input modes. By varying the temperature of one PLCU fiber, we tune its optical length and record the resulting four‑fold coincidence counts $n_{\mathrm{Corr}}$ between the two beam-splitter outputs and their corresponding heralds, correcting for higher photon‑number contributions as described in the Supplementary Material. This configuration effectively realizes Hong–Ou–Mandel (HOM) interference between distinct sources [58, 59, 60]. Given the measured spectro-temporal mode number $K=1.05$ from Figure 3 b, the expected maximum visibility is $V\approx 95\%$. Some measurements approach this limit, while slight reductions in others are attributed primarily to imperfections in specific 50:50 beam splitters implemented in the interferometer.
@@
Violation of Classicality Bounds
@@ Discussion
In the GBS configuration, however, increasing the brightness of the input SMSV states does not yield a stronger or more statistically significant violation of the classical bound beyond a certain limit (Figure 4 dark orange line). The maximum violation exceeds $26\,\sigma$ at an average photon number of $\langle n\rangle=0.172$, while the local minimum at $\langle n\rangle=0.569$ corresponds to a $14\,\sigma$ violation. Once the mean photon number exceeds the value at this local minimum, the system’s quantumness no longer continues to increase. Furthermore, Figure 4 reveals that at a mean photon number of $\langle n\rangle=2.152$, the two-mode correlators from the GBS data do not exhibit quantumness for over 90% of the input phase values. Therefore, a classical model could efficiently reproduce this data.
Understanding the cause for the observed reduction in quantumness is paramount. One natural hypothesis is that the reduced quantumness results from increased input-state impurity caused by system losses, consistent with recent work that decomposes impure SMSV states into a classical thermal component and a pure squeezed component [25]. Simulations performed using The Walrus [62] library, however, reveal that this behavior is compatible with simulations based on a lossless system (see Supplementary Material), asking for explanations beyond input state impurity. Furthermore, the simulations were run with fixed input phases for each run and were calculated up to a photon number of 15—also ruling out the possibility that the reduction in quantumness stems from phase drifts or limited PNR.
@@
Methods
Pump Source
Pump pulses are generated by a newly developed Menlo laser system that begins with a SmartComb system producing femtosecond pulses in the telecom C‑band. Light from the SmartComb enters a laser extension unit that provides optical gain and generates various control signals. Finally, the amplified pulses are frequency‑doubled in a free‑space second‑harmonic stage to yield picosecond pulses centered around 772 nm with a full-width at half maximum bandwidth of approximately 1 nm, a repetition rate of 80 MHz and pulse energies up to approximately 12 nJ.
[…]
Squeezed‑State Generation
The prepared pump‑pulse train enters the squeezed‑light source, which comprises a waveguided parametric down‑conversion (PDC) source, a temporal compensation crystal (CC), EOM2, a PBS, and fiber‑coupling optics. The 20 mm‑long waveguided PPKTP crystal (AdvR Inc.) is designed for a type‑II PDC process that creates signal and idler fields that are spectrally indistinguishable and that generates ps-long squeezed light pulses in a single spectro-temporal mode.
Following the waveguide, the output passes through a filtering stage composed of a silicon window that suppresses the pump and a 2 nm spectral filter that removes phase matching side lobes. An 8.8 mm‑long KTP group‑delay compensation crystal (CC) then corrects for birefringent walk‑off accumulated in the waveguide, ensuring temporal overlap of the signal and idler pulses. The filtered, temporally synchronized fields subsequently enter an electro‑optic modulator (EOM2, QUBIG GmbH HVOS‑SWIR‑1.5k100), which sets the output polarization of the transmitted modes on a polarizing beam splitter (PBS). The PBS output ports are fiber‑coupled. Interference of the two polarized fields at the PBS produces independent SMSV states used for GBS implementations, whereas deterministic separation of the fields yields a TMSV state used to realize SBS or TBS configurations in a single experimental data run.
Multiplexed Detector and Heralding
A multiplexed detection module attached to one PBS output both spatially and temporally separates the incoming pulses to measure the photon number in this heralding arm. This module enables photon‑number heralding for implementing higher-order SBS and allows verification of the lack of correlations between the two PBS outputs in the GBS configuration. Discarding the heralding component of the data obtained SBS effectively prepares thermal input states for the interferometer [57].
Time‑to‑Space Demultiplexing
Light exiting the second PBS output passes through an active time‑to‑space demultiplexer (DMX16_80, QUBIG GmbH), which routes each input pulse into individual output fibers via an EOM switching tree with an efficiency exceeding 80%. Each output fiber is fed into a temperature‑stabilized path‑length compensation unit (PLCU, Menlo Systems GmbH) consisting of segments of fiber of variable length. These include approximately 60 m of fiber wound around a piezo actuator and additional lengths compensating the 12.5 ns time difference between consecutive pulses. The optical path length can be tuned using temperature control and mechanical strain.
Integrated Interferometer
Photon‑Number Detection
Photon numbers in the eight utilized output modes of the interferometer are measured using eight SNSPDs (Single Quantum) with quantum efficiency exceeding 90% and timing jitter below 20 ps. An intrinsic PNR scheme is implemented, relying on the $\approx 2$ ps pulse duration of the squeezed‑light states and the sub‑2 ps jitter of the Swabian Instruments Time Tagger X. Using this technique, our implementation discriminates events with up to three photons per pulse with high confidence [55, 63].

### v2

@@ Abstract
The original boson sampling paradigm—consisting of multiple single-photon input states, a large interferometer, and multi‑channel click detection—was originally proposed as a photonic route to quantum computational advantage. Its non‑Gaussian resources, essential for outperforming any classical system, are provided by single‑photon inputs and click detection. Yet the drive toward larger experiments has led to the replacement of experimentally demanding single‑photon sources with Gaussian states, thereby diminishing the available non‑Gaussianity—a critical quantum resource. As the community broadens its focus from the initial sampling task to possible real‑world applications, it becomes crucial to quantify the performance cost associated with reducing non‑Gaussian resources and to benchmark sampling platforms that employ different input states. To address this need, we introduce the Paderborn Quantum Sampler (PaQS), a hybrid platform capable of performing sampling experiments with eight Gaussian or non‑Gaussian input states in a 12‑mode interferometer within a single experimental run. This architecture enables direct, side‑by‑side benchmarking of distinct sampling regimes under otherwise identical conditions. By employing a semi-device-independent framework, offering certification that does not rely on prior knowledge of the interferometer or the input states, we verify that the observed data cannot be reproduced using classical resource states—a prerequisite for demonstrating quantum advantage. Applying this framework, we observe clear performance gains arising from non‑Gaussian input states.
preprint: APS/123-QED
I Introduction
@@
The non‑Gaussian element in GBS systems arises from photon‑number‑resolved (PNR) detection (or in earlier implementations, click detection). Extending the computational capabilities of such systems is therefore most easily achieved by supplementing this resource, either through advanced operations within the interferometer [31, 34, 35, 36], or by driving the interferometer with more exotic non‑Gaussian input states, whose impact has been the subject of extensive investigation [37, 38, 39, 40, 41, 42, 43, 44, 45]. Consequently, the next generation of sampling devices will require a conceptual redesign of both their system architecture and resource allocation.
@@
A corresponding rethink is necessary for benchmarking the performance of these newly conceived sampling systems. A commonly used verification technique is to quantify the closeness of generated data to that which is expected from theory, typically implemented using a metric known as the total variation distance [1]. However, this method requires exponentially many samples and, therefore, does not scale. Furthermore, the total variation distance is sensitive to experimental imperfections; errors in implementing the multi-mode interferometer will strongly affect the total variation distance. Progressing the field therefore requires new certification tools capable of identifying genuinely non‑classical features—or quantumness—and experimental platforms that enable fair, side‑by‑side comparison of different architectures. These measures should provide some degree of insensitivity to experimental imperfections and should be accessible from the generated data.
@@
In this work, we introduce the Paderborn Quantum Sampler (PaQS), an experimental platform designed to benchmark multiple sampling schemes within a single experimental run. To compare the performance associated with different input‑state resources—both Gaussian and non-Gaussian—as illustrated in Figure 1, we further develop a benchmarking framework based on normally ordered moments of the photon‑number operator [46, 47, 48, 49, 50, 51]. This framework, in contrast to previous approaches, certifies the presence of quantumness in generated data – a quantity that has been proven as a necessity for obtaining complexity in sampling experiments [52]. Furthermore, this approach is robust to experimental imperfections and can be directly applied to data generated from the boson sampling platform — unlike measures such as Wigner function negativity [53], which require additional, complex measurement apparatus.
@@
Although one might expect, a priori, that Fock states would outperform squeezed states due to their higher degree of nonclassicality [54], the situation is more nuanced: it has been shown that the specific type of nonclassicality that can be exploited for a given task is not necessarily correlated with the “absolute” amount of nonclassicality [55, 56]. Therefore, understanding the benefits offered by distinct input states is critical: in fact, our results reveal pronounced differences between the generated GBS and SBS datasets. The quantumness of SBS data is seen to steadily increase with the mean photon number of the input states, whereas the GBS data displays strong non‑classical signatures at low mean photon numbers, but fails to maintain them as the brightness increases – even though it is known that the nonclassicality of squeezed states increases with increasing degree of squeezing [54]. This divergence underscores the fundamentally different behaviors arising from distinct input‑state resources and highlights the need for a conceptual shift when evaluating the performance of boson sampling architectures.
@@
Figure 1: Conceptual implementations for various sampling configurations. In all cases, the input fields enter from the bottom of the setup and are detected with PNR detectors upon exiting the interferometer at the top. The input states of the interferometer can be either eight SMSV states, heralded Fock states, or thermal states, corresponding to implementing GBS, SBS, and thermal boson sampling (TBS), respectively. To implement SBS, the second mode of each TMSV input state is detected with PNR detectors to herald the number of photons entering each interferometer port. To realize sampling with thermal states the second mode of each TMSV state is discarded, i.e., traced out.
II Experimental Setup
@@
Figure 2 presents a schematic of the PaQS system, divided into its constituent subsystems for clarity. The design of PaQS emphasizes integration: both the squeezed-light source (or parametric down-conversion source, PDC) and the programmable interferometer are realized on integrated photonic platforms. Notably, the integrated source, in conjunction with an electro-optic modulator and polarizing beamsplitter, enables dynamic switching between TMSV- and SMSV-state generation. This architecture allows multiple sampling configurations to be acquired during a single experimental run, thereby ensuring that results from the different sampling schemes are directly commensurable. Furthermore, the picosecond squeezed-light pulses generated by the waveguided source enables the implementation of intrinsic photon-number resolution (PNR) in the employed superconducting nanowire single-photon detectors (SNSPDs). The integrated interferometer offers convenient and complete programmability of the implemented unitary transformation, with an average insertion loss below 3 dB. An overview of the key subsystems of PaQS is given in Figure 2 and described here, while detailed descriptions can be found in the Appendixes.
II.1 Pump Source
@@
Pump pulses are generated by a newly developed laser system that begins with a SmartComb (Menlo Systems GmbH) system producing femtosecond pulses in the telecom C‑band. Light from the SmartComb enters a laser extension unit that provides optical gain and generates various control signals. The amplified pulses are then frequency‑doubled in a free‑space second‑harmonic stage to yield picosecond pulses centered around 772 nm with a full-width at half maximum bandwidth of approximately 1 nm, a repetition rate of 80 MHz and pulse energies up to approximately 12 nJ.
[…]
II.2 Squeezed-State Generation
The prepared pump‑pulse train enters the 20 mm‑long waveguided PPKTP crystal (AdvR Inc.) designed for a type‑II PDC process that creates spectrally indistinguishable, ps-long signal and idler fields in a single spectro-temporal mode. After the waveguide crystal, the output fields pass through a filtering stage composed of a silicon window that suppresses the pump and a 2 nm spectral filter that removes phase matching side lobes. An 8.8 mm‑long KTP group‑delay compensation crystal (CC) then corrects for birefringent walk‑off accumulated in the waveguide, ensuring temporal overlap of the signal and idler pulses. The filtered, temporally synchronized fields subsequently enter an electro‑optic modulator (EOM2, QUBIG GmbH HVOS‑SWIR‑1.5k100), which sets the output polarization of the transmitted modes on a polarizing beam splitter (PBS). By varying the voltage applied to EOM2, we can rapidly switch (rise/fall time of 4 ns) between generating two independent SMSV states and a TMSV state at the PBS output. Interference of the two polarized fields at the PBS produces independent SMSV states used for GBS implementations [8], whereas deterministic separation of the fields yields a TMSV state used to realize SBS [7] or TBS [58] configurations in a single experimental data run. The fields at both PBS output ports are fiber‑coupled.
II.3 Multiplexed Detector and Heralding
A multiplexed detection module attached to one PBS output both spatially and temporally separates the incoming pulses to measure the photon number in this heralding arm. This module enables photon‑number heralding for implementing higher-order SBS and allows verification of the lack of correlations between the two PBS outputs in the GBS configuration. Discarding the heralding component of the data obtained in SBS effectively prepares thermal input states for the interferometer [57].
II.4 Time‑to‑Space Demultiplexing
The fiber-coupled light that exited the second PBS output passes through an active time‑to‑space demultiplexer (DMX16_80, QUBIG GmbH), which routes each input pulse into individual output fibers via an EOM switching tree with an efficiency exceeding 80%. Each output fiber is fed into a temperature‑stabilized path‑length compensation unit (PLCU, Menlo Systems GmbH) consisting of segments of fiber of variable length. These include approximately 60 m of fiber wound around a piezo actuator and additional lengths compensating the 12.5 ns time difference between consecutive pulses. The optical path length can be tuned using temperature control and mechanical strain.
II.5 Integrated Interferometer
II.6 Photon‑Number Detection
Photon numbers in the eight utilized output modes of the interferometer are measured using eight SNSPDs (Single Quantum) with quantum efficiency exceeding 90% and timing jitter below 20 ps. An intrinsic PNR scheme is implemented, relying on the $\approx 2$ ps pulse duration of the squeezed‑light states and the sub‑2 ps jitter of the Swabian Instruments Time Tagger X [59, 60]. Using this technique, our implementation discriminates events with up to three photons per pulse with high confidence [60, 61].
III System Characterization
@@
We summarize below the principal methods used to characterize and verify the operation of the PaQS subsystems. The corresponding results are shown in Figure 3, with further details provided in Appendix B.
@@
Single spectro-temporal-mode operation of our type-II waveguide source is verified through second-order autocorrelation measurements [57]. The system operates in SBS mode while photon statistics are recorded for one output of the PBS as the mean photon number is varied by adjusting the pump power driving the nonlinear process. As shown in Figure 3 b, the second-order correlation function converges to $g^{(2)}=1.95\pm 0.03$ at higher power, corresponding to an expected Schmidt mode number $K=1.05\pm 0.03$ effective modes—confirming both spectro-temporal purity and suitability for sampling experiments. The increase in $g^{(2)}$ at low mean photon numbers originates from a residual SMSV contribution because of imperfect polarization separation at the PBS (see Appendix B).
Next, we assess the ability of PaQS to switch between SMSV and TMSV generation. Operating in the low‑gain regime, we record the coincidence counts $n_{\mathrm{C}}$ between the herald and signal detectors as the driving voltage of EOM2’s high-voltage amplifier is varied (Figure 3 c). The normalized coincidence signal varies nearly sinusoidally with a fitted visibility of $V=96.3\%\!\pm\!1.2\%$ (See Appendix B), demonstrating precise and continuous tunability between SBS (maximum coincidences) and GBS (minimum coincidences) configurations. The difference in number of normalized coincidence counts for the two displayed SBS configurations (at 0 V and $\approx$1.35 V) is due to the different spatial-mode profiles of the two polarization modes within the waveguide.
Finally, Figure 3 d illustrates the precision with which we can tune the optical path-length difference between input modes using the PLCUs in order to optimize the indistinguishability between different sources (further details in Appendix B). Implementing the SBS configuration in the low‑gain regime, we set the interferometer to realize a 50:50 beam splitter between any pair of chosen input modes. By varying the temperature of one PLCU fiber, we tune its optical length and record the resulting four‑fold coincidence counts $n_{\mathrm{Corr}}$ between the two beam-splitter outputs and their corresponding heralds, correcting for higher photon‑number contributions as described in Appendix B. This configuration effectively realizes Hong–Ou–Mandel (HOM) interference between distinct sources [63, 64, 65]. Given the measured spectro-temporal mode number $K=1.05$ from Figure 3 b, the expected maximum visibility is $V\approx 95\%$. Some measurements approach this limit, while slight reductions in others are attributed primarily to imperfections in specific 50:50 beam splitters implemented in the interferometer.
@@
IV Violation of Classicality Bounds
@@ V Discussion
In the GBS configuration, however, increasing the brightness of the input SMSV states does not yield a stronger or more statistically significant violation of the classical bound beyond a certain limit (Figure 4 dark orange line). The maximum violation exceeds $26\,\sigma$ at an average photon number of $\langle n\rangle=0.172$, while the local minimum at $\langle n\rangle=0.569$ corresponds to a $14\,\sigma$ violation. Once the mean photon number exceeds the value at this local minimum, the system’s quantumness no longer continues to increase. Furthermore, Figure 4 reveals that at a mean photon number of $\langle n\rangle=2.152$, the two-mode correlators from the GBS data do not exhibit quantumness for over 90% of the input phase values. Therefore, using classical resources alone, one could reproduce this data.
Understanding the cause for the observed reduction in quantumness is paramount. One natural hypothesis is that the reduced quantumness results from increased input-state impurity caused by system losses, consistent with recent work that decomposes impure SMSV states into a classical thermal component and a pure squeezed component [25]. Simulations performed using The Walrus [68] library, however, reveal that this behavior is compatible with simulations based on a lossless system (see Appendix D), asking for explanations beyond input state impurity. Furthermore, the simulations were run with fixed input phases for each run and were calculated up to a photon number of 15—also ruling out the possibility that the reduction in quantumness stems from phase drifts or limited PNR.
@@
Acknowledgements.
M.S would like to acknowledge many fruitful discussions with various colleagues and friends: Sevag Gharibian, Dhruva Sambrani, Hamid Naeij, Tina Chen, Dorian Rudolph, Jonas Kamminga, Robert Schade, Nathan Walk, Helen Chrzanowski, Jens Eisert, Dario Cilluffo, Martin Plenio, Andrew White, Fatemeh Mohit, Craig Hamilton, Shivani Singh, Aurèl Gàbris, Magdalena Paryzkova, Igor Jex, Tim Ralph, Ryan Marshman, Nibedita Swain. F.L would like to thank Laura Serino for help with the SLM. F.S. is part of the Max Planck School of Photonics supported by the Dieter Schwarz Foundation, the German Federal Ministry of Research, Technology and Space (BMFTR), and the Max Planck Society. T.S and T.B acknowledge partial funding by the European Union (ERC, QuESADILLA, 101042399). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them. This work has received funding from the German Federal Ministry of Research, Technology and Space within the PhoQuant project (Grant No. 13N16103).
Author contributions
The research project was initiated and directed by C.S., who oversaw its conceptual development and execution. M.S. and K.-H.L. designed the experiment with input from B.B, and C.S.; K.-H.L. and S.A. built the experimental setup with the help of F.L., J-L.E., C.P., F.Sc., J.L., M.S., M.Y., A.K., C.S., M.L., O.K., and F.St.; M.Y., A.K., and R.H. developed and/or partially installed the laser system and path length compensation units; F.Sc. designed and constructed the multiplexed detection scheme for heralding with aid from J.L.; M.Y., S.A., K.-H.L. implemented the PLCU scheme; F.Sc., T.S., T.B., J.L. developed and implemented the PNR detection scheme; F.L. implemented the pulse shaper with help from K.-H.L., and F.Sc.; K.-H.L. designed and built the squeezed-light source with the help of F.L., J-L.E., and F.Sc.; J.-L.E., K.-H.L., S.A, M.R. performed the experimental runs and data collection with the help of J.L. and C.P.; J.-L.E. completed data analysis with the help of L.A., J.S., J.L., and C.P.; M.S. prepared the manuscript with the help of K.-H.L., S.A., J-L.E., F.L., C.P., F.Sc., L.A., J.S., M.R., B.B., C.S., and J.L.; All authors discussed the results and reviewed the manuscript.